Shift report 20000629 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Thu Jun 29 2000 - 23:59:03 EDT

  • Next message: http account: "Shift report 20000630 00:00-08:00"

    ------------------------------------------------------------
    Date: Thu Jun 29 23:59:02 2000, Shift: 16:00-24:00
    Supervisor: FR, with: FJ, ZIS
    Log book#2 pages 95 - 96
    
    Shift Summary:
    ------------------------------------------------------------
     - No beam during the entire shift. There was a problem
       with Booster main magnet PS. An attempt to ramp to
       gamma=70 started at 22:40 but the ramp failed.
       Current status "Ramping back down to injection".
    
     - KO has changed the C1 ADC mapping (see Log book p.95).
       Run 1570 is a pedestal run taken with the new mapping.
    
     - Problems : 1) Middle NIM bin in Rack 3 was found off. 
       We turned it on. 2) Lower NIM bin in Rack 4 was found
       again with pulser GG not working. It recovered after
       cycling power. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Fri Jun 30 2000 - 00:03:10 EDT